# Semantic Mapping in Indoor Embodied AI – A Comprehensive Survey and Future Directions

### v1

@@ 3.2 What is the structure of this map?
A semantic map can be structured as a spatial grid map, topological map, point-cloud map or a hybrid map (Fig. 4). A spatial grid map is a top-down grid where each grid cell represents an area in the physical environment. So if an object is at a certain location $(X,Y,Z)$ in a 3D scene, the semantic map will contain information about that object at the corresponding grid cell $(x,y)$, where $x$ and $y$ are the row and column numbers respectively such that there is a direct mapping from $(X,Y,Z)$ to $(x,y)$. For navigation, most spatial maps are 2D, such that a grid cell ignores the $Z$-axis (up direction) by aggregating the semantic information across the up axis. However they can also be 3D where the $Z$-axis is divided into discrete bins. On the other hand, a topological map is a graph-like structure where nodes represent objects or important landmarks in the scene and edges represent relationship (distance, spatial relation, etc.) between them. It is also possible to store semantic information on a point-cloud map, which can be viewed as a 3D map with varying density. In a point-cloud map, information is associated with each point $(x,y,z)$ corresponding to 3D location $(X,Y,Z)$ in the physical space. Unlike the voxel-grid, which is regularly spaced, points can be sampled at varying densities. Some works combine two or more of the above structures to form a hybrid map since each structure has its own advantages and limitations. We discuss each of these in detail in Sec. 4.
3.3 What is stored in this map?
@@ 3.4 How is the map built?
3.5 Considerations in map building
Despite the huge progress in map building in the embodied AI community, there are many challenging aspects that need continual research and improvement.
Real-time processing. When building the map online, the agent needs to continually process sensory data and update the map in real-time by accumulating semantic information. This is a computationally intensive task and hence requires efficient algorithms. Moreover, in dynamic environments such as autonomous driving, the algorithms need to have low latency.
Memory consumption and scalability. Creating and maintaining large-scale semantic maps require sufficient storage and may be difficult to scale. Such large maps may result from navigating a large environment such as outdoor cityscapes or navigating intricate indoor spaces for a long period of time or storing dense semantic features inside the map. Efficiently updating such large-scale maps is still a challenge that remains to be solved (Gu et al., 2023).
Noise and uncertainty. In real-world robots the data from sensors (cameras, IMU, etc.) might be noisy due to reflective surfaces, uneven gradient, etc. Such noisy data introduces uncertainty in the map built by the robot. However, in the embodied AI community, it is a common practice to assume noiseless sensors in simulated environments, which results in a noiseless map. When transferred directly from simulation to real-world, the map building methods might not work very well, unless special techniques to deal with noise and uncertainty are employed (Chaplot et al., 2020a; Georgakis et al., 2022b).
Dynamic environments. In many applications such as autonomous driving or robotic surgery, the environment is always changing because of moving objects, changing lighting conditions and evolving structures. The map building techniques need to be adaptable to these dynamic environmental changes and build accurate and consistent map.
Semantic understanding. Identifying objects in an environment accurately is crucial for creating useful semantic maps. This relies on advanced computer vision techniques which have seen a huge progress from identifying a fixed set of objects (He et al., 2017; Ren et al., 2015) to identifying ‘any’ object in an image (Zhou et al., 2022; Liu et al., 2023b). Moreover, contextual understanding and spatial relationships between objects is useful for effective decision-making in complex semantic tasks (Antol et al., 2015; Gordon et al., 2018).
Usability and reliability. Creating interactive interfaces for humans to query semantic maps is a new area many researchers are focusing on (Peng et al., 2023; Yamazaki et al., 2023). Such queryable maps have not yet been used in any of the downstream embodied AI tasks, but remains a promising direction to explore. Moreover, ensuring that the built maps are reliable and consistent induces trust in real-world safety-critical applications. This requires building accurate maps of the environment through rigorous validation and testing.
Standardization. Developing standardized frameworks for map building might help in collaboration and integration across different systems and platforms. Although the robotics community has standardized practices in map building (e.g. SLAM), the embodied AI community relies on different techniques (Gupta et al., 2017; Chaplot et al., 2020a) and may benefit from a common framework.
@@ 4 Map structure
In this section we will look at various map structures that have been used in prior works (see Tab. 2). A semantic map can be structured in various ways: spatial grid map, topological map, point-cloud map or a hybrid map. Spatial grid maps are metric maps of the environment such that its dimensions align to that of the environment and they can be structured as either 2D or 3D grids. Topological maps, on the other hand, represent the environment through a set of landmarks represented as nodes and relation between adjacent landmarks represented as edges in the form of a graph. Point-cloud maps are the densest form of 3D maps whose all three dimensions align to the 3D space such that each 3D point in the scene is captured in the map. Two or more of these types of maps are sometimes combined together to form Hybrid maps.
@@ 4.1 Spatial grid map
Table 3: Spatial grid maps. Prior works build spatial grid maps for various embodied AI tasks. Information is aggregated onto the map over time in many different ways such as learned recurrent networks and replacing with most recent information among others.
@@ 4.2 Topological map
Compared to the high-precision grid maps, topological maps are graph-like structures with nodes connected to each other by edges. This essentially abstracts a large space into significant areas (nodes) where the agent can take decisions and connections or paths between them (edges) (Johnson, 2018). This enables parsing the environment into a local and a global structure such that the agent can plan locally in the small space represented as nodes while navigating the large space through graph search following the edges. This way of planning and navigating is inspired from how humans navigate in an unseen environment in that they identify and memorize significant landmarks and find paths to reach those landmarks (Janzen & Van Turennout, 2004; Foo et al., 2005; Chan et al., 2012; Epstein & Vass, 2014). Thus topological maps have been a popular choice in both traditional robotics research (Thrun & Montemerlo, 2006; Lorbach et al., 2014; Rosinol et al., 2020b; Campos et al., 2021) as well as in embodied AI research (Savinov et al., 2018; Chen et al., 2021; Chaplot et al., 2020b; Kwon et al., 2021; Gu et al., 2023; Mehan et al., 2024; Garg et al., 2024; An et al., 2024; Yang et al., 2024).
@@
Table 4: Topological maps. Various works in indoor embodied AI build topological map either in an exploration phase or online while performing the task. The nodes often store learned features about the observation or temporal information (visitation timestep), while edges may store relative poses between a pair of nodes or types of edges.
@@
Method Map Building Phase Node values Node Feature Encoder Edge values
SPTM (2018) Pre-Exploration Image features ResNet18
@@
NTS (2020b) Online Image features ResNet18 Relative pose in polar coordinates
@@
CMTP (2021) Pre-Exploration Image features ResNet152 Relative pose in discrete polar coordinates (8 directions, 3 distances (0-2m,2-5m,>5m)
VGM (2021) Online Image features, visitation timestep ResNet18
@@
TSGM (2023) Online ImageNode stores image features; ObjectNode stores features for detected objects pretrained image and object encoders
@@
RoboHop (2024) Online Image features for each image segment CLIP (2021), DINOv2 (2023) edge types denoting inter- and intra-image connectivity
@@
A key design decision during topological map building is what should be represented as nodes and what should be edges. Generally speaking, the nodes encode semantic information about locations in the environment such that the agent can make a decision whereas the edges store relationship or connection between the nodes. For indoor navigation, the landmarks for the nodes are typically objects in the environments. They can also be openings or intersections (Fredriksson et al., 2023), locations the agents has visited (Chaplot et al., 2020b), and other regions of interest (Kim et al., 2023; Shah et al., 2023; Garg et al., 2024). For navigation, two nodes are connected with an edge if it is possible to navigate from one node to another. Some methods also store spatial relationships between the nodes (Gu et al., 2023) in the edges to enable better reasoning.
@@
Another important aspect in the topological map creation is how to determine if two observations are similar to each other, in which case the two are mapped to the same node. If they are not similar, two different nodes exist for the two observations. This requires the map building methods to compare RGB images. The goal here is to classify two images as similar if (1) they are exactly the same or (2) there is a slight change in direction or distance between the two. Traditionally this is the problem of data association in SLAM-based systems, where an incoming observation could be matched to multiple landmarks (nodes) and either the best match is selected or a new node is created to mark a new landmark (Bowman et al., 2017; Dellaert et al., 2000). In embodied AI some works use a pretrained classifier network to implement this. The network is trained to classify whether two images are from the same area. NTS uses MLP trained with a cross-entropy loss in a supervised manner to predict whether are similar. This however needs annotated pairs of training data. Cross-Modal Transformer Planner (CMTP) (Chen et al., 2021), on the other hand, uses an oracle ‘Reachability Estimator’ to first obtain the geodesic distance between the two underlying locations based on the traversibility of the 3D mesh. If the distance is below a threshold, it maps them to the same node. Visual Graph Memory (VGM) (Kwon et al., 2021) also uses a pretrained network to determine if two images are similar. But they learn an unsupervised representation of the observations which are then projected onto an embedding space. The idea is to have the embeddings of observations coming from nearby areas clustered together because they are likely to have similar appearances. The training data in this case consist of randomly sampled observations from the training environments, thus eliminating the need for manual annotations. Kim et al. (2023) on the other hand use semantic similarity score obtained from a pretrained network (Li et al., 2021) between two images to determine whether they are the same nodes. A similar approach is taken in LM-Nav (Shah et al., 2023) where they use CLIP to calculate the cosine similarity between image features. Tab. 4 compares different methods that build topological maps.
@@
4.3 Point-cloud map
It is also possible to accumulate semantic information directly on to the 3D geometry of a scene, either triangle meshes or point clouds (which are increasingly popular as they are easier to work with than triangle meshes). For instance, given a point cloud, we can associate semantic information with each point in the scene. While not traditionally considered maps, we can view these representations as a type of semantic map as they associate semantics with spatial information. Often, a neural network is used to obtain the features associated with each point given the position (x,y,z) as input. Such neural functions that transform coordinates to real-valued vectors are also known as neural fields. These neural fields can produce dense 3D semantic map where each 3D point in the scene is captured and represented in the map.
Although traditionally such representations are mostly built for 3D scene understanding tasks, there has been increasing interest in using them for embodied AI. In scene understanding tasks, the goal is to perform various semantic inferences for each 3D point (Peng et al., 2023; Xu et al., 2024). In other words, given a point-cloud of a scene, the goal is to perform semantic segmentation, affordance estimation, room type classification, 3D object search and so on. Representations learned for such tasks are also often useful downstream for embodied AI tasks.

### v2

@@ 3.2 What is the structure of this map?
A semantic map can be structured as a spatial grid map, topological map, dense geometric or a hybrid map (Fig. 1). A spatial grid map is a top-down grid where each grid cell represents an area in the physical environment. So if an object is at a certain location $(X,Y,Z)$ in a 3D scene, the semantic map will contain information about that object at the corresponding grid cell $(x,y)$, where $x$ and $y$ are the row and column numbers respectively such that there is a direct mapping from $(X,Y,Z)$ to $(x,y)$. For navigation, most spatial maps are 2D, such that a grid cell ignores the $Z$-axis (up direction) by aggregating the semantic information across the up axis. However they can also be 3D where the $Z$-axis is divided into discrete bins. On the other hand, a topological map is a graph-like structure where nodes represent objects or important landmarks in the scene and edges represent relationship (distance, spatial relation, etc.) between them. It is also possible to store semantic information on a dense geometric, which can be viewed as a 3D map with varying density. In a dense geometric, information is associated with each point $(x,y,z)$ corresponding to 3D location $(X,Y,Z)$ in the physical space. Unlike the voxel-grid, which is regularly spaced, points can be sampled at varying densities. Some works combine two or more of the above structures to form a hybrid map since each structure has its own advantages and limitations. We discuss each of these in detail in Sec. 4.
3.3 What encoding is stored in this map?
@@ 4 Map structure
In this section we will look at various map structures that have been used in prior works (see Tab. 1) in more depth. A semantic map can be structured in various ways: spatial grid map, topological map, dense geometric or a hybrid map. Spatial grid maps are metric maps of the environment such that its dimensions align to that of the environment and they can be structured as either 2D or 3D grids. Topological maps, on the other hand, represent the environment through a set of landmarks represented as nodes and relation between adjacent landmarks represented as edges in the form of a graph. Dense geometric maps are the densest form of 3D maps whose all three dimensions align to the 3D space such that each 3D point in the scene is captured in the map. Two or more of these types of maps are sometimes combined together to form Hybrid maps.
@@ 4.1 Spatial grid map
Table 3: Spatial grid maps. Prior works build spatial grid maps (2D or 3D) for various embodied tasks. Information is aggregated onto the map over time in many different ways such as learned recurrent networks and replacing with most recent information among others.
@@
Efficient (2013) 3D scene analysis robot VMR-Oakland (2011), Freiburg (2012) most recent
@@
L2M (2021) ObjectNav Habitat (2019) MP3D (2017), ObjectNav (2020b) most recent
SEAL (2021) ObjectNav Habitat (2019) custom with Gibson (2018) scenes channel-wise max-pool
CM2 (2022c) Vision and Language Nav Habitat (2019) VLN-CE (2020) most recent
RNR-Map (2023) Visual Nav Habitat (2019) Gibson (2018) mean
Le-RNR-Map (2023) Visual Nav Habitat (2019) Gibson (2018) mean
VoxPoser (2023b) Table-top manipulation SAPIEN (2020), robot custom -
@@
Instruction-guided (2025) Vision-language navigation Habitat VLN-CE (2020) max-pool
@@ 4.2 Topological map
Compared to the high-precision grid maps, topological maps are graph-like structures with nodes connected to each other by edges. This essentially abstracts a large space into significant areas (nodes) where the agent can take decisions and connections or paths between them (edges) (Johnson, 2018). This enables parsing the environment into a local and a global structure such that the agent can plan locally in the small space represented as nodes while navigating the large space through graph search following the edges. This way of planning and navigating is inspired from how humans navigate in an unseen environment in that they identify and memorize significant landmarks and find paths to reach those landmarks (Janzen & Van Turennout, 2004; Foo et al., 2005; Chan et al., 2012; Epstein & Vass, 2014). Moreover, in instruction-following tasks, where a robot follows instructions (e.g., ‘go to the yellow chair in front of you’), the language instruction is often represented as a graph with high initial uncertainty. As the robot moves and observes its environment, it uses a probabilistic model to update this graph, leading to improved task performance (Tucker et al., 2019; Patki et al., 2019; Arkin et al., 2020a; Raychaudhuri et al., 2025). Thus topological maps have been a popular choice in traditional robotics research such as SLAM (Thrun & Montemerlo, 2006; Rosinol et al., 2020b; Campos et al., 2021), manipulation (Arkin et al., 2020b; Patki et al., 2020), object search (Aydemir et al., 2011; 2013; Lorbach et al., 2014), language grounding (MacMahon et al., 2006; Chen & Mooney, 2011b; Tellex et al., 2011a; Matuszek et al., 2013; Duvallet et al., 2013; Gong & Zhang, 2018; Paul et al., 2018) as well as in embodied AI research (Savinov et al., 2018; Chen et al., 2021; Chaplot et al., 2020b; Kwon et al., 2021; Gu et al., 2023b; Mehan et al., 2024; Garg et al., 2024; An et al., 2024; Yang et al., 2024; Tang et al., 2025).
@@
Table 4: Topological maps. Various works in indoor embodied AI build topological map either in an exploration phase or online while performing the task. The nodes often store explicit, learned features about the observation or temporal information (visitation timestep), while edges may store relative poses between a pair of nodes or types of edges.
@@
Method Task Map Building Phase Node values Node Feature Encoder Edge values
Topo SLAM (2001) SLAM Online Pose
SLAM++ (2013) SLAM Online Pose, object type
Imitation (2013) Instruction following Online Pose
TopoMap (2018) SLAM Online free space
SPTM (2018) ImageNav Pre-Exploration Image features ResNet18 (2016)
Compact (2019) Instruction following Online Semantic attributes
Multimodal (2020a) Instruction following Online Semantic attributes
@@
NTS (2020b) ImageNav Online Image features ResNet18 Relative pose in polar coordinates
@@
CMTP (2021) Instruction following Pre-Exploration Image features ResNet152 Relative pose in discrete polar coordinates (8 directions, 3 distances (0-2m,2-5m,>5m)
VGM (2021) ImageNav Online Image features, visitation timestep ResNet18
Fast (2022) Path planning Pre-Exploration free space
@@
TSGM (2023) ImageNav Online ImageNode stores image features; ObjectNode stores features for detected objects pretrained image and object encoders
LM-Nav (2023) Instruction following Online Image features CLIP (2021)
@@
RoboHop (2024) Language querying Online Image features for each image segment CLIP, DINOv2 (2023) edge types denoting inter- and intra-image connectivity
Clio (2024) Mobile manipulation Online object semantics CLIP
LIFGIF (2025) Instruction following Online Semantic attributes
ASHiTA (2025) Task planning Online Task, sub-task, objects MobileCLIP (2024)
@@
A key design decision during topological map building is what should be represented as nodes and what should be edges. Generally speaking, the nodes encode semantic information about locations in the environment such that the agent can make a decision whereas the edges store relationship or connection between the nodes. For indoor navigation, the landmarks for the nodes are typically objects in the environments. They can also be openings or intersections (Fredriksson et al., 2023), locations the agents has visited (Chaplot et al., 2020b), and other regions of interest (Kim et al., 2023; Shah et al., 2023; Garg et al., 2024). For navigation, two nodes are connected with an edge if it is possible to navigate from one node to another. Some methods also store spatial relationships between the nodes (Gu et al., 2023b) in the edges to enable better reasoning.
@@
Another important aspect in the topological map creation is how to determine if two observations are similar to each other, in which case the two are mapped to the same node. If they are not similar, two different nodes exist for the two observations. This requires the map building methods to compare RGB images. The goal here is to classify two images as similar if (1) they are exactly the same or (2) there is a slight change in direction or distance between the two. Traditionally this is the problem of data association in SLAM-based systems, where an incoming observation could be matched to multiple landmarks (nodes) and either the best match is selected or a new node is created to mark a new landmark (Bowman et al., 2017; Dellaert et al., 2000). In embodied AI some works use a pretrained classifier network to implement this. The network is trained to classify whether two images are from the same area. NTS uses MLP trained with a cross-entropy loss in a supervised manner to predict whether are similar. This however needs annotated pairs of training data. Cross-Modal Transformer Planner (CMTP) (Chen et al., 2021), on the other hand, uses an oracle ‘Reachability Estimator’ to first obtain the geodesic distance between the two underlying locations based on the traversibility of the 3D mesh. If the distance is below a threshold, it maps them to the same node. Visual Graph Memory (VGM) (Kwon et al., 2021) also uses a pretrained network to determine if two images are similar. But they learn an unsupervised representation of the observations which are then projected onto an embedding space. The idea is to have the embeddings of observations coming from nearby areas clustered together because they are likely to have similar appearances. The training data in this case consist of randomly sampled observations from the training environments, thus eliminating the need for manual annotations. Kim et al. (2023) on the other hand use semantic similarity score obtained from a pretrained network (Li et al., 2021a) between two images to determine whether they are the same nodes. A similar approach is taken in LM-Nav (Shah et al., 2023) where they use CLIP to calculate the cosine similarity between image features. Tab. 4 compares different methods that build topological maps.
@@
4.3 Dense geometric map
Semantic information can also be directly accumulated onto the 3D geometry of a scene using triangle meshes (Valentin et al., 2013), surfels (Stückler & Behnke, 2014) or point clouds (Jatavallabhula et al., 2023), with point cloud maps being the most widely used due to simplicity and ease of use. While these representations are not always traditionally considered “maps”, they effectively function as semantic maps by coupling geometric structure with semantic content. Dense geometric map representations, particularly point clouds, have long been widely used in robotics for tasks such as mapping, localization and navigation, as well as in broader 3D scene understanding tasks (Peng et al., 2023; Xu et al., 2024). Recently, there has been a surge of interest in applying these representations to embodied AI for downstream reasoning and decision-making (Gu et al., 2023b). We further categorize dense geometric map representations into point cloud maps (Sec. 4.3.1) and neural fields (Sec. 4.3.2). Point cloud maps store 3D points with associated semantic labels whereas neural fields represent scenes as continuous functions using neural networks (see Tab. 5 for a summary of prior works).
Table 5: Dense geometric maps allow storing semantics onto the 3D geometry of a scene. It can be further categorized into point-cloud maps and neural fields. Here, VL is visual-and-language, Str is Structure, PC is point-cloud, SF is surfel, GS is gaussian splat and NF is neural fields.
. Method Task Env Str Encoding Train 6D-SLAM (2007) SLAM robot PC geometry Towards semantic maps (2008) SLAM robot PC 3D location, category label, robot pose Robust 3D mapping (2009) SLAM robot PC geometry 3D object-class map (2012) SLAM robot SF category label Parsing (2012) Semantic segmentation robot PC 3D location, category label Street Scenes (2012) Semantic segmentation simulator PC 3D location, category label Dense 3D (2014) 3D reconstruction simulator PC 3D location, category label SemanticFusion (2017) SLAM robot SF 3D location, normal, category label Droid-SLAM (2021) SLAM robot PC geometry Voldor+ SLAM (2021) SLAM robot PC geometry Semantic-NeRF (2021) Scene understanding Habitat, Replica NF category label NeSF (2021) Semantic segmentation Kubric NF category label 3D aware ObjNav (2022a) Object navigation Habitat PC 3D location, semantic category, consistency information ConceptFusion (2023) Rearrangement, autonomous driving, semantic segmentation Replica, robot PC 3D location, normal, confidence, color, pixel aligned VL features OpenScene (2023) Open-vocab scene understanding ScanNet, MP3D, nuscenes PC CLIP LERF (2023) Open-vocab scene understanding robot NF learned neural VL features LGM (2023) Open-vocab mobile manipulation robot GS CLIP dense patch features CLIP-Fields (2023) Object navigation Habitat, robot NF learned neural VL features LangSplat (2024) Open-vocab semantic segmentation LERF GS VL features from CLIP, SAM SemanticGaussians (2024) Open-vocab scene understanding ScanNet, LERF GS VL features from CLIP, SAM, LSeg GaussNav (2024) Visual navigation Habitat GS category label GeFF (2024) Open-vocab mobile manipulation robot GS MaskCLIP (2022) SGS-SLAM (2024a) SLAM robot GS category label GaussianGrasper (2024b) Open-vocab tabletop manipulation robot GS CLIP GeomGS (2025) Localization robot GS geometric confidence
4.3.1 Point cloud maps
Point clouds have become increasingly popular over the years due to their ease of use. Given a 3D point cloud, semantic labels can be associated with each point and thus by coupling geometric structure with semantic content, they serve as a dense semantic map. It is also possible to spatially extend each point to a 3D Gaussian with a covariance matrix. Such representations are called Gaussian splats, and allow for differentiable rendering. Point clouds (and Gaussian splats more recently) have been adopted in a variety of robotics tasks as well as in 3D scene understanding tasks, gaining popularity in the embodied AI community in recent years.
In robotics, point cloud-based mapping has undergone significant evolution, driven by advances in sensing, computational power and SLAM algorithms. Early efforts (Nüchter et al., 2007; May et al., 2009) use 3D sensor scans to build geometrically consistent representations of large-scale environments by emphasizing accuracy and loop-closure. As real-time visual SLAM has matured, dense point cloud generation from RGB-D sensors became feasible, enabling detailed indoor reconstructions (Whelan et al., 2015; Dai et al., 2017; Schops et al., 2019; Min & Dunn, 2021). More recently, learning-based SLAM systems like Droid-SLAM (Teed & Deng, 2021) integrate deep feature extraction and end-to-end optimization to improve SLAM. CoFusion (Rünz & Agapito, 2017) introduce systems for dense SLAM with semantic fusion, pushing toward maps that are not only geometrically accurate but also semantically meaningful. This trend has been accelerated recently with ConceptFusion (Jatavallabhula et al., 2023) and Clio (Maggio et al., 2024), which align open-vocabulary visual features (e.g., CLIP) with 3D reconstructions to create task-agnostic and queryable maps. These efforts represent a shift from purely geometric maps toward dense, semantically-rich, open-vocabulary representations that support a wide range of downstream tasks. Moreover, Gaussian splatting has recently emerged as a powerful alternative to traditional point cloud representations in robotics, enabling high-fidelity scene reconstruction and differentiable 3D representations useful for tasks such as visual localization, SLAM, and manipulation. Recent works such as SGS-SLAM (Li et al., 2024a) demonstrate how to include semantic labels onto Gaussian splats to effectively perform SLAM, and GaussianGrasper (Zheng et al., 2024b) demonstrates how these representations can be fused with semantics to support open-vocabulary grasping.
In embodied AI, maintaining point cloud maps typically prove to be memory-intensive and not ideal for real-time use. However some efforts have been made to improve efficiency. For instance, Zhang et al. (2023a) builds a 3D semantic scene representation based on an online point cloud-based construction algorithm (Zhang et al., 2020), made efficient by using a tree-based dynamic data structure. This method is very memory efficient and a lot faster when applied to the ObjectNav task. Recent efforts also use Gaussian splatting in navigation tasks. For example, Lei et al. (2024), based on Gaussian splatting, achieves state-of-the-art performance on the ImageNav task.
In 3D scene understanding, point cloud representations have become a cornerstone due to their flexibility and fine spatial resolution. Early learning-based approaches such as PointNet (Qi et al., 2017a) and PointNet++ (Qi et al., 2017b) enable direct processing of unordered point sets for tasks like semantic segmentation and object classification. These were followed by more expressive models like KPConv (Thomas et al., 2019), which introduce learnable convolutions, and Point Transformer (Zhao et al., 2021), which leverage attention mechanisms to capture long-range dependencies across points. More recent work integrates multi-modal and language supervision into point cloud understanding(Peng et al., 2023; Xu et al., 2024), thus enabling fine-grained inference such as semantic segmentation, affordance prediction and 3D object search. OpenScene (Peng et al., 2023), for example, predicts dense 3D features so that they are co-embedded with the corresponding text and the image in the CLIP embedding space. This allows the association of each 3D point in the scene with semantic information such that the scene can be queried using text to infer physical properties, affordances, etc. CLIP2Scene (Chen et al., 2023c) also uses CLIP to perform a 3D point cloud segmentation on outdoor scenes for application in autonomous driving. However, because the optimization happens per scene in many of these methods, they are typically expensive and not suitable for real-time use in embodied applications.
Summary. Point cloud maps remain a widely used representation in the research community due to their simplicity, high fidelity and direct correspondence with sensor data. However, challenges persist in handling noise, achieving scalability in large environments and maintaining consistent semantics over time. Future research should aim to integrate multi-modal features and improve robustness to dynamic scenes in point cloud maps, all the while maintaining efficiency.
4.3.2 Neural fields